# Towards Optimal Circuit Generation: Multi-Agent Collaboration Meets Collective Intelligence

### v1

@@ Abstract
Large language models (LLMs) have transformed code generation, yet their application in hardware design produces gate counts 38%–1075% higher than human designs. We present CircuitMind, a multi-agent framework that achieves human-competitive efficiency through three key innovations: syntax locking (constraining generation to basic logic gates), retrieval-augmented generation (enabling knowledge-driven design), and dual-reward optimization (balancing correctness with efficiency). To evaluate our approach, we introduce TC-Bench, the first gate-level benchmark harnessing collective intelligence from the TuringComplete ecosystem—a competitive circuit design platform with hundreds of thousands of players. Experiments show CircuitMind enables 55.6% of model implementations to match or exceed top-tier human experts in composite efficiency metrics. Most remarkably, our framework elevates the 14B Phi-4 model to outperform both GPT-4o mini and Gemini 2.0 Flash, achieving efficiency comparable to the top 25% of human experts without requiring specialized training. These innovations establish a new paradigm for hardware optimization where collaborative AI systems leverage collective human expertise to achieve optimal circuit designs. Our model, data, and code are open-source at [link will be released upon acceptance].
@@ I Introduction
TABLE I: Comparison of LLM-Generated and Human-Optimized Circuits
@@
Synthesis via Yosys 0.34 [1] (LLM: Gemini 2.0 Flash, Human: Top 25%)
@@
This evolution mirrors compiler development, which progressed from simple translation to sophisticated multi-pass optimization. However, a critical insight remains overlooked: hardware design requires specialized reasoning processes that a single-agent LLM fundamentally cannot replicate. The solution to the Boolean Optimization Barrier lies not in creating larger models or more domain-specific datasets, but in rethinking the fundamental architecture of AI-based hardware design systems.
@@
  • Human-Competitive Results: CircuitMind enables 55.6% of implementations to match or exceed top-tier human experts, with our Phi-4 implementation outperforming both GPT-4o mini and Gemini 2.0 Flash.
@@ II-A Gate-Level Circuit Optimization
where $\alpha,\beta$ are weighting parameters. This NP-hard problem requires non-local reasoning. Traditional logic synthesis tools like Yosys and ABC employ heuristic minimization algorithms and technology mapping to tackle this complexity, while human experts apply Boolean algebra manipulation, Karnaugh maps, shared logic identification, and technology mapping across entire circuit structures.
@@ II-B Measuring Design Efficiency
where $G$ represents gate count and $D$ represents delay. We use $\alpha=\beta=1$ for balanced evaluation, reflecting common optimization goals in TuringComplete. However, real-world circuit optimization often involves complex trade-offs between area, delay, and power, and the TuringComplete scoring mechanism might incorporate additional factors like simulation ticks. For benchmark-level aggregation, we use geometric mean:
@@ II-C3 Phase 3: Early Physical Optimization (2024-present)
Recent approaches like BetterV [8] and CodeV [9] incorporate physical considerations but remain limited by generating behavioral code that synthesis tools must convert to gates. As Table II shows, while frameworks like RTLSquad use multi-agent systems for RTL optimization and GenEDA focuses on reasoning about netlists, no existing framework enforces direct gate-level reasoning or aligns with human expert optimization practices in the same manner as CircuitMind. The qualitative assessments in Table 2 reflect the primary focus of each framework based on their published methodologies; for instance, BetterV uses a discriminator for optimization guidance, while HiVeGen targets PPA but still operates at the RTL level.
@@
RTLSquad Multi-Agent (RTL Opt.) No High No
GenEDA Encoder-Decoder (Netlist Reasoning) No Low No
@@ II-D The Boolean Optimization Barrier
LLMs generate sequences based on statistical patterns rather than performing the structured symbolic reasoning required for Boolean optimization. Human experts apply systematic minimization techniques like Quine-McCluskey algorithms that LLMs cannot effectively replicate.
@@ II-E Benchmark Misalignment
A key challenge in advancing LLM-based hardware design is the disconnect between evaluation metrics and physical implementation quality. As shown in Table III, existing benchmarks have evolved from functional correctness (VerilogEval [4]) to limited PPA metrics (RTLLM [12]) and more comprehensive quality assessments (ResBench[13], GenBen[14]), but most lack alignment with human expert performance. TC-Bench addresses this gap by leveraging collective intelligence from TuringComplete’s competitive environment to establish meaningful gate-level optimization targets based on actual human performance tiers.
Despite these advancements, TC-Bench offers a unique contribution by harnessing collective intelligence from a competitive gaming environment (TuringComplete) to establish gate-level optimization challenges and human performance tiers based on the SEI metric. This approach provides a distinct perspective focused on direct gate-level optimization efficiency derived from a large community actively competing on such tasks. Table III provides an updated comparison.
@@
TC-Bench Gate Func. Corr., SEI Game (TuringComplete) Yes (Tiers) Intelligence source, SEI metric
VerilogEval Verilog (RTL) Func. Corr. Academic/Web Limited Focus on functional correctness
RTLLM RTL Func. Corr., PPA Academic/Web Limited PPA metrics, reference designs
ResBench Verilog (RTL) Func. Corr., FPGA Synthetic/Academic No FPGA resource efficiency
GenBen Mixed Func. Corr., QoR, Debug Diverse (Silicon+) Limited QoR focus, contamination handling
RTL-Repo Multi-file RTL Func. Corr. GitHub Limited Focus on multi-file projects
@@ II-F The Fundamental Reasoning Gap
Existing approaches attempt to bridge the efficiency gap through three primary strategies, all of which face significant limitations. Scale-based approaches fail because the barrier is fundamentally algorithmic rather than capacity-related; simply increasing model size does not address the underlying reasoning limitations in Boolean optimization. Fine-tuning strategies are severely hampered by the extreme scarcity of optimized gate-level examples in available training data, creating an insurmountable data bottleneck that prevents effective specialization. Single-agent verification loops, while improving functional correctness, lack the specialized reasoning capabilities necessary for effective Boolean optimization and cannot replicate the structured symbolic manipulation that human experts employ.
@@ III TC-Bench: Harnessing Collective Intelligence
To address the benchmark misalignment problem identified earlier, we introduce TC-Bench—a gate-level evaluation framework that leverages collective intelligence from competitive circuit designers on the TuringComplete platform [^1] https://turingcomplete.game/. This framework establishes meaningful efficiency targets for AI-generated circuits based on human performance benchmarks.
@@ III-A Collective Intelligence Approach
TC-Bench derives its data from TuringComplete, where a large community competes to optimize digital circuits using minimal logic gates while meeting functional requirements. Solutions are ranked based on gate count and performance (delay), generating thousands of expert-refined implementations that often significantly exceed the efficiency of conventional synthesis tool outputs. TC-Bench harvests these optimized designs across player skill levels (Figure 2(a)), creating a multi-tiered reference framework that provides diverse, human-optimized implementations rather than relying on standard synthesis results.
Figure 2: TC-Bench’s collective intelligence harvesting process: (a) Data extraction from TuringComplete leaderboards across skill tiers. (b) Benchmark stratification into Easy, Medium, and Hard challenges.
@@ III-B Benchmark Structure
TC-Bench comprises 28 challenges selected based on circuit complexity, functional diversity, and optimization potential, stratified into three difficulty levels (Figure 2(b)):
@@ III-C1 Human Reference Tiers
  • Top tier (Rank 1-250): Expert optimizers. SEI range: [0.0905, 0.1252].
  • Mid tier (Rank 250-750): Skilled practitioners. SEI range: [0.0856, 0.0905).
  • Low tier (Rank 750-1000): Competent designers. SEI range: [0.0851, 0.0856).
@@ IV-B2 Retrieval-Augmented Generation
When receiving a design task, CoderAgent first classifies the circuit type (combinational or sequential) and identifies required building blocks. It then queries the knowledge database for optimized implementations of these components, incorporating them directly into generation prompts as concrete examples of efficient design patterns. This approach enables the model to leverage past successes while adapting to new requirements, creating a bootstrapping mechanism that progressively enhances design capabilities [16].
@@ IV-C Collaborative Workflow
Executor performs functional verification through compilation and simulation, with results fed back to Reviewer for potential optimization.
@@ V-B Comparative Performance Analysis
The results, detailed in Table IV and summarized in Figure 5, reveal key patterns analyzed below.
@@ V-B1 Functional Correctness and Physical Efficiency Enhancements
CircuitMind delivers substantial improvements in both functional correctness and physical efficiency across all evaluated models. For functional correctness (Figure 5, top), the most striking gains appear on hard tasks, where baseline LLMs typically struggled. Notable improvements include DeepSeek-R1 (0.26 to 0.99, 3.8×3.8$\times$3.8 × increase) and Phi-4 (0.02 to 0.26, 13×13$\times$13 × increase) on hard tasks.
For physical efficiency (Figure 5, bottom), CircuitMind consistently improves overall SEI for all tested models. The efficiency gains are particularly significant on hard tasks, with some models showing order-of-magnitude improvements: CM(GPT-4o mini) achieved 90×90$\times$90 × higher SEI and CM(Phi-4) 89×89$\times$89 × higher SEI compared to their respective baselines. While some models showed slight SEI decreases on easy tasks, the substantial improvements on medium and hard tasks resulted in significant overall efficiency gains, demonstrating CircuitMind’s effectiveness in complex optimization scenarios.
@@ V-B2 Model-Specific Performance Patterns
Figure 6 confirms these observations, showing that high-performing implementations like CM(Gemini-2.0) and CM(DeepSeek-R1/V3) successfully solve problems across all difficulty categories, with particularly strong performance on hard tasks that baseline models struggled to address.
@@ V-C Human-Competitive Performance Evaluation
To assess CircuitMind’s real-world significance, we compared its performance against the human expert tiers established in TC-Bench (Figure LABEL:fig:model_size_sei).
@@ V-D Component Contribution Analysis
To isolate the impact of our key technical innovations, we conducted an ablation study focused on the Retrieval-Augmented Generation (RAG) components. We compared three configurations, all maintaining Syntax Locking and Dual-Reward mechanisms:
@@ V-E Case Study: Boolean Optimization
VI Discussion and Conclusion
Our work addresses two critical challenges in LLM-based circuit design: the fundamental reasoning limitations of language models in Boolean optimization and the lack of human-aligned evaluation metrics. We introduce CircuitMind, a multi-agent framework, and TC-Bench, a benchmark based on collective intelligence, to overcome these challenges.
VI-A Key Innovations and Implications
The integration of Syntax Locking with Retrieval-Augmented Generation represents a significant advancement in AI-based hardware design. This approach enables models to perform sophisticated Boolean optimization without requiring specialized gate-level training data. Our results demonstrate that architectural innovation can effectively overcome the fundamental reasoning limitations inherent in individual LLMs.
The empirical validation is compelling: 55.6% of CircuitMind implementations match or exceed top-tier human experts in physical efficiency metrics. Most notably, the Phi-4-based implementation achieves top-tier human-competitive results through the strategic decomposition of complex design tasks across specialized agents. This success demonstrates that collaborative reasoning can unlock capabilities that individual models cannot achieve independently.
Simultaneously, TC-Bench addresses a critical evaluation gap in the field by providing the first standardized framework that measures both functional correctness and physical efficiency against human expert references. By establishing realistic efficiency targets derived from competitive human designers, TC-Bench aligns model development with practical design objectives that matter in real-world applications.
VI-B Limitations and Future Work
While CircuitMind demonstrates significant advances in gate-level optimization, several important limitations remain. First, the current implementation focuses primarily on Boolean optimization and delay reduction without explicitly addressing power consumption or other industrial-relevant physical constraints. Second, the knowledge evolution mechanism relies on examples from a relatively small benchmark set, which would require more sophisticated management techniques for scaling to complex industrial designs.
TC-Bench, while valuable for standardized evaluation, does not fully capture industrial-scale circuit design complexity. Real-world design challenges involve additional considerations such as power management, timing closure, and manufacturability constraints that are not yet represented in our benchmark.
Future research should extend both the framework and benchmark to address these limitations. This includes developing more comprehensive metrics that incorporate power efficiency, expanding the knowledge base to handle larger-scale designs, and enhancing the collaborative mechanisms to tackle more complex optimization scenarios with industrial relevance.
VI-C Conclusion
CircuitMind demonstrates that multi-agent collaboration guided by human-aligned evaluation metrics can significantly narrow the gap between AI-generated and human-optimized circuit designs. Our research shows that architectural innovation in reasoning distribution offers a more promising path than simply scaling model size or expanding training data.
The complementary innovations of syntax-constrained generation and retrieval-augmented optimization enable human-competitive efficiency without requiring extensive gate-level training data. This approach has implications beyond hardware design, potentially serving as a model for similar collaborative frameworks in domains requiring specialized reasoning and physical constraint satisfaction.
By establishing TC-Bench as a human-aligned benchmark, we contribute a valuable resource for future hardware design automation research. Our work represents a promising direction for bridging the gap between AI capabilities and human expertise across complex engineering domains where physical efficiency and functional correctness must be balanced.

### v2

@@ Abstract
Large language models (LLMs) have transformed code generation, yet their application in hardware design produces gate counts 38%–1075% higher than human designs. We present CircuitMind, a multi-agent framework that achieves human-competitive efficiency through three key innovations: syntax locking (constraining generation to basic logic gates), retrieval-augmented generation (enabling knowledge-driven design), and dual-reward optimization (balancing correctness with efficiency). To evaluate our approach, we introduce TC-Bench, the first gate-level benchmark harnessing collective intelligence from the TuringComplete ecosystem—a competitive circuit design platform with hundreds of thousands of players. Experiments show CircuitMind enables 55.6% of model implementations to match or exceed top-tier human experts in composite efficiency metrics. Notably, our framework elevates the 14B-parameter Phi-4 model (ranked 6th among 1,000 human experts on TuringComplete) to outperform both GPT-4o mini and Gemini 2.0 Flash without requiring specialized training. These innovations establish a new paradigm for hardware optimization where collaborative AI systems leverage collective human expertise to achieve optimal circuit designs. Our model, data, and code are open-source at [link will be released upon acceptance].
@@ I Introduction
TABLE I: Comparison of LLM-Generated and Human-Optimized Circuits in TuringComplete[1]
@@
Synthesis via Yosys[2] (LLM: Gemini 2.0 Flash)
@@
This evolution mirrors compiler development, progressing from simple translation to sophisticated multi-pass optimization. However, a critical insight remains overlooked: hardware design requires specialized reasoning processes that a single-agent LLM fundamentally cannot replicate. The solution to the Boolean Optimization Barrier lies not in larger models or more domain-specific datasets, but in rethinking the fundamental architecture of AI-based hardware design systems.
@@
  • Human-Competitive Results: CircuitMind enables 55.6% of implementations to match or exceed top-tier human experts, with our Phi-4[3] implementation outperforming both GPT-4o mini[4] and Gemini 2.0 Flash.
@@ II-A Gate-Level Circuit Optimization
where $\alpha,\beta$ are weighting parameters. This NP-hard problem requires non-local reasoning. Traditional logic synthesis tools like Yosys[2] and ABC [5] employ heuristic minimization algorithms and technology mapping to tackle this complexity, while human experts apply Boolean algebra manipulation, Karnaugh maps, shared logic identification, and technology mapping across entire circuit structures.
@@ II-B Measuring Design Efficiency
where $G$ represents gate count and $D$ represents delay. We use $\alpha=\beta=1$ for balanced evaluation, reflecting common optimization goals in TuringComplete [1]. However, real-world circuit optimization often involves complex trade-offs between area, delay, and power, and the TuringComplete [1] scoring mechanism might incorporate additional factors like simulation ticks. For benchmark-level aggregation, we use geometric mean:
@@ II-C3 Phase 3: Early Physical Optimization (2024-present)
Recent approaches like BetterV [12] and CodeV [13] incorporate physical considerations but remain limited by generating behavioral code that synthesis tools must convert to gates. As Table II shows, while frameworks like RTLSquad [14] use multi-agent systems for RTL optimization and GenEDA [15] focuses on reasoning about netlists, no existing framework enforces direct gate-level reasoning or aligns with human expert optimization practices in the same manner as CircuitMind. The qualitative assessments in Table 2 reflect the primary focus of each framework; for instance, BetterV[12] uses a discriminator for optimization guidance, while HiVeGen [16] targets PPA but still operates at the RTL level.
@@
RTLSquad [14] Multi-Agent (RTL Opt.) No High No
GenEDA [15] Encoder-Decoder (Netlist Reasoning) No Low No
@@ II-D The Boolean Optimization Barrier
LLMs generate sequences based on statistical patterns rather than performing the structured symbolic reasoning required for Boolean optimization. Human experts apply systematic minimization techniques like Quine-McCluskey algorithms[18] that LLMs cannot effectively replicate.
@@ II-E Benchmark Misalignment
A key challenge in advancing LLM-based hardware design is the disconnect between evaluation metrics and physical implementation quality. As shown in Table III, existing benchmarks have evolved from functional correctness (VerilogEval [8]) to limited PPA metrics (RTLLM [19]) and more comprehensive quality assessments (ResBench[20], GenBen[21]), but most lack alignment with human expert performance. TC-Bench addresses this gap by leveraging collective intelligence from TuringComplete’s[1] competitive environment to establish meaningful gate-level optimization targets based on actual human performance tiers.
Despite these advancements, TC-Bench offers a unique contribution by harnessing collective intelligence from a competitive gaming environment to establish gate-level optimization challenges and human performance tiers based on the SEI metric. This approach provides a distinct perspective focused on direct gate-level optimization efficiency derived from a large community actively competing on such tasks. Table III provides an updated comparison.
@@
TC-Bench(ours) Gate Func. Corr., SEI Game (TuringComplete) Yes (Tiers) Intelligence source, SEI metric
VerilogEval[8] RTL Func. Corr. Academic/Web Limited Focus on functional correctness
RTLLM[19] RTL Func. Corr., PPA Academic/Web Limited PPA metrics, reference designs
ResBench[20] RTL Func. Corr., FPGA Synthetic/Academic No FPGA resource efficiency
GenBen[21] Mixed Func. Corr., QoR, Debug Diverse (Silicon+) Limited QoR focus, contamination handling
RTL-Repo[22] Multi-file RTL Func. Corr. GitHub Limited Focus on multi-file projects
@@ II-F The Fundamental Reasoning Gap
Existing approaches attempt to bridge the efficiency gap through three primary strategies, all facing significant limitations. Scale-based approaches fail because the barrier is fundamentally algorithmic rather than capacity-related; simply increasing model size does not address the underlying reasoning limitations in Boolean optimization. Fine-tuning strategies are hampered by extreme scarcity of optimized gate-level examples, creating an insurmountable data bottleneck. Single-agent verification loops, while improving functional correctness, lack the specialized reasoning capabilities necessary for effective Boolean optimization and cannot replicate the structured symbolic manipulation that human experts employ.
@@ III TC-Bench: Harnessing Collective Intelligence
To address the benchmark misalignment problem identified earlier, we introduce TC-Bench—a gate-level evaluation framework that leverages collective intelligence from competitive circuit designers on the TuringComplete platform. This framework establishes meaningful efficiency targets for AI-generated circuits based on human performance benchmarks.
@@ III-A Collective Intelligence Approach
TC-Bench derives its data from TuringComplete, where a large community competes to optimize digital circuits using minimal logic gates while meeting functional requirements. Solutions are ranked based on gate count and performance (delay), generating thousands of expert-refined implementations that often significantly exceed the efficiency of conventional synthesis tool outputs. TC-Bench harvests these optimized designs across player skill levels (Fig. 2(a)), creating a multi-tiered reference framework that provides diverse, human-optimized implementations rather than relying on standard synthesis results.
Figure 2: TC-Bench’s collective intelligence harvesting process
@@ III-B Benchmark Structure
TC-Bench comprises 28 challenges selected based on circuit complexity, functional diversity, and optimization potential, stratified into three difficulty levels (Fig. 2(b)):
@@ III-C1 Human Reference Tiers
  • Top tier (Rank 1-10): Expert optimizers. SEI range: [0.0951, 0.1252].
  • Mid tier (Rank 11-300): Skilled practitioners. SEI range: [0.0905, 0.0924].
  • Low tier (Rank 301-1000): Competent designers. SEI range: [0.0851, 0.0905].
@@ IV-B2 Retrieval-Augmented Generation
When receiving a design task, CoderAgent first classifies the circuit type (combinational or sequential) and identifies required building blocks. It then queries the knowledge database for optimized implementations of these components, incorporating them directly into generation prompts as concrete examples of efficient design patterns. This approach enables the model to leverage past successes while adapting to new requirements, creating a bootstrapping mechanism that progressively enhances design capabilities.
@@ IV-C Collaborative Workflow
Executor performs functional verification through compilation and simulation, with results fed back to Reviewer for potential optimization. This dynamic prompting process between CoderAgent and Reviewer is illustrated in Fig. 4.
@@ V-A Experimental Setup
We evaluated CircuitMind using the Pass@k metric from [28], defined as:
[…]
where $k$ is samples per task, $n$ is total samples, and $c$ is correct solutions. With $n=20$, we estimated correctness from correct sample fractions. Pass@k reflects the probability of solving a task with the best of $k$ samples.
@@ V-B Comparative Performance Analysis
The results, detailed in Table IV and summarized in Fig. 5, reveal key patterns analyzed below.
@@ V-B1 Functional Correctness and Physical Efficiency Enhancements
CircuitMind delivers substantial improvements in both functional correctness and physical efficiency across all evaluated models. For functional correctness (Fig. 5, top), the most striking gains appear on hard tasks, where baseline LLMs typically struggled. Notable improvements include DeepSeek-R1 (0.26 to 0.99, 3.8×3.8$\times$3.8 × increase) and Phi-4 (0.02 to 0.26, 13×13$\times$13 × increase) on hard tasks.
For physical efficiency (Fig. 5, bottom), CircuitMind consistently improves overall SEI for all tested models. The efficiency gains are particularly significant on hard tasks, with some models showing order-of-magnitude improvements: CM(GPT-4o mini) achieved 90×90$\times$90 × higher SEI and CM(Phi-4) 89×89$\times$89 × higher SEI compared to their respective baselines. While some models showed slight SEI decreases on easy tasks, the substantial improvements on medium and hard tasks resulted in significant overall efficiency gains, demonstrating CircuitMind’s effectiveness in complex optimization scenarios.
@@ V-B2 Model-Specific Performance Patterns
Fig. 6 confirms these observations, showing that high-performing implementations like CM(Gemini-2.0) and CM(DeepSeek-R1/V3) successfully solve problems across all difficulty categories, with particularly strong performance on hard tasks that baseline models struggled to address.
@@ V-C Human-Competitive Performance Evaluation
To assess CircuitMind’s real-world significance, we compared its performance against the human expert tiers established in TC-Bench (Fig. 7).
Figure 7: SEI vs. Model Size Compared to Human Expert Tiers.
@@ V-D Component Contribution Analysis
To isolate the impact of our key technical innovations, we conducted an ablation study focused on the RAG components. We compared three configurations, all maintaining Syntax Locking and Dual-Reward mechanisms:
@@ V-E Case Study: Boolean Optimization
VI Conclusion
This research addressed the critical efficiency gap in LLM-based hardware design, identifying the Boolean Optimization Barrier and benchmark misalignment as key obstacles. Our proposed multi-agent framework, CircuitMind, integrated with the human-aligned evaluation capabilities of TC-Bench, demonstrated a significant advancement in tackling these challenges.
Experimental results clearly show that our approach substantially improves both functional correctness and, crucially, the physical efficiency (SEI) of AI-generated circuits across diverse LLMs. Achieving human-competitive performance in 55.6% of implementations, notably with a smaller model like Phi-4 outperforming larger ones, underscores a core finding: intelligent collaborative architecture and targeted optimization strategies, as embodied in CircuitMind, are more pivotal than raw model scale for overcoming gate-level reasoning limitations. TC-Bench proved essential in quantifying this progress against realistic human expert benchmarks.
While current work focuses on gate count and delay, limitations remain regarding power optimization, industrial constraints, and knowledge base scalability. Future efforts should aim to incorporate these broader physical design considerations and enhance the collaborative mechanisms for complex, real-world scenarios.
In conclusion, this work validates that structuring AI systems for collaborative, specialized reasoning, guided by benchmarks reflecting human expertise, offers a powerful pathway to achieving human-level performance in complex engineering optimization tasks like digital circuit design. This paradigm shift emphasizes architectural innovation and holds promise for advancing AI-driven automation in hardware and potentially other physically constrained domains.